8. Elasticity problems:
a. The world demand for crude oil is estimated to have a short-run price elasticity of 0.05. If the initial price of oil were $100 per barrel, what would be the effect on oil price and quantity of an embargo that curbed world oil supply by 5 percent? (For this problem, assume that the oil-supply curve is completely inelastic.)
b. To show that elasticities are independent of units, refer to Table 3-1. Calculate the elasticities between
each demand pair. Change the price units from dollars to pennies; change the quantity units from millions of boxes to tons, using the conversion factor of 10,000 boxes to 1 ton. Then recalculate the elasticities in the fi rst two rows. Explain why you get the same answer.
1
Expert's answer
2021-10-05T11:11:23-0400

Elasticity of demand indicates the percentage change in quantity demanded due to one percent change in price.

a.

The supply of the oil is assumed to be inelastic the supply curve is assumed to be vertical. Considering the scenario, the oil supply would reduce by 5 per cent shifting it to the left.



There will be an increase in price of oil and decline in the quantity demanded. The percentage change in quantity demanded is 5 percent, and there is a reduction in supply.

Calculate the change in price as follows:

If E(D), the elasticity of demand, is 0.05, then:

"E(D)=\frac{percentage\; change\;in\;quantity\;demanded}{percentage\;change\;in\;price} \\

0.05=\frac{5}{percentage\;change\;in\;price} \\"

Percentage change in price = "\frac{5}{0.05}=100"


Thus, the price of oil will be

$30x2=$60 per barrel

Therefore, there will be a 100 percent increase price from$30 to $60 per barrel, and there will be a decline in quantity by 5 percent.

b.



If the quantities are changed: multiply the p’s by 100 (100 pennies in a dollar), and Q’s by 100.

Calculate as follows:



The elasticities remain same. It means that the elasticities are independent of unit. The reason is that the elasticity is calculated using the percentage changes on both the numerator and the denominator.